Circular No. 7771

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ie IN UGC 5542 A. Dimai, Cortina d’Ampezzo, Italy, reports the discovery by A. Bin-

coletto, San Donà di Piave, Italy, of an apparent supernova (mag ∼ 16.5) on an image taken on Dec. 9.13 UT with a 0.30-m f/5 reflector (limiting mag ∼ 17.5). SN 2001ie is located at α = 10^{h}16^{m}50^{s}.70, δ = +60^{o}16′44′′.5 (equinox 2000.0), which is ∼ 15′′ west and 15′′ south of the nucleus of UGC 5542. A confirming image taken on Dec. 10.02 with the 0.5-m telescope at Col Drusci Observatory shows the new object at mag 16.1. An image taken by Bincoletto on Jan. 21 showed nothing at the position of the new object (limiting mag ∼ 17.5), and it does not appear on Palomar Sky Survey red and blue plates.

SUPERNOVA 2001if NEAR MCG +06-6-43 G. M. Hurst, Basingstoke, England, reports the discovery by T. Boles,

Coddenham, England, of an apparent supernova (mag 17.2) on a 30-s unfiltered CCD image taken on Dec. 9.986 UT in the course of the U.K. Nova/Supernova Patrol with the 0.36-m ‘A’ Schmidt-Cassegrain reflector. The new object, which appeared at similar brightness on images taken on Dec. 10.016 and 10.788, is located at α = 2^{h}28^{m}54^{s}.93, δ = +38^{o}05′52′′.7 (equinox 2000.0; average from the three images), which is 70′′ west and 3′′ south of the nucleus of MCG +06-6-43. M. Armstrong, Rolvenden, England, reports that a CCD image taken on Dec. 10.785 also shows the new star. Boles notes that SN 2001if is absent from his master image taken on Sept. 10.997 (limiting mag 18.5) and was not recorded on Palomar Sky Survey images taken in 1986 (red, limiting mag 21) and 1993 (blue, limiting mag 20). Hurst adds that MCG +06-6-43 is the brightest galaxy in the discovery field, but there are several other galaxies nearby and the parent galaxy is not entirely certain.

CATALOGUE OF COMETARY ORBITS 2001 A new edition of the Catalogue of Cometary Orbits, with entries com-

plete through November, is now available. The 152-page Catalogue sells for $40.00 ($60.00 for airmail delivery outside North America). There is also an e-mail edition (with just the basic orbital information) for $75.00. Orders should be sent to Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A., by FAX to +1-617-495-7231 or by e-mail to mlohmiller@cfa.harvard.edu. Payment can be by credit card or by check made payable to Smithsonian Astrophysical Observatory.

2001 December 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green